OpenStudy (princesssleelee):

The function g is defined as follows for the domain given. g(x)=1-3x, domain = {-1,0,1,2} write the range of g using set notation. Then graph g.

OpenStudy (princesssleelee):

@agent0smith

OpenStudy (princesssleelee):

@Mehek14

OpenStudy (kgrendel0324):

I think its all real numbers

OpenStudy (princesssleelee):

That isnt an option, its a fill in the blank for range=____

OpenStudy (kgrendel0324):

oh... I'm not sure then.

OpenStudy (agent0smith):

g(x)=1-3x, domain = {-1,0,1,2} the domain is your x values. plug them in, one at a time, and that will give you the y-values.

OpenStudy (princesssleelee):

I dont understand

OpenStudy (agent0smith):

For example, using the equation g(x)=1-3x, plug in the first x value which is -1: 1-3(-1) = 4 so 4 is one of the values in the range. Now plug in 0, then the other two.